# Recursive construction and enumeration of self-orthogonal and self-dual codes over finite commutative chain rings of even characteristic - I

### v1

@@ 4 Enumeration formulae for self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}$
In this section, we will obtain explicit enumeration formulae for self-orthogonal and self-dual codes of an arbitrary length over $\mathscr{R}_{e,m}.$ In particular, when $e=2,$ one can easily see that either $\mathscr{R}_{e,m}\simeq\mathbb{F}_{2^{m}}[u]/\langle u^{2}\rangle$ or $\mathscr{R}_{e,m}\simeq GR(4,m).$ The enumeration formula for self-dual codes over $\mathbb{F}_{2^{m}}[u]/\langle u^{2}\rangle$ was derived by Gaborit [14, Th. 3], while the enumeration formula for self-orthogonal codes over $\mathbb{F}_{2^{m}}[u]/\langle u^{2}\rangle$ was obtained by Galvez et al. [15, Th. 2]. Yadav and Sharma [31, Th. 5.1] counted all self-orthogonal and self-dual codes of an arbitrary length over $GR(4,m).$ In view of this, we assume, throughout this section, that $e\geq 3.$ We also assume that $n$ is a positive integer and $\lambda_{1},\lambda_{2},\ldots,\lambda_{e+1}$ are non-negative integers satisfying $n=\lambda_{1}+\lambda_{2}+\cdots+\lambda_{e+1}.$ Now, let $\mathfrak{S}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})$ and $\mathfrak{U}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})$ denote the number of distinct self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m},$ respectively. We note, by Lemma 2.1 and Remark 2.1, that $\mathfrak{S}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})=0$ if $2\lambda_{1}+2\lambda_{2}+\cdots+2\lambda_{e-j+1}+\lambda_{e-j+2}+\cdots+\lambda_{j}>n$ holds for some integer $j$ satisfying $s+1\leq j\leq e,$ while $\mathfrak{U}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})=0$ if $\lambda_{j}\neq\lambda_{e-j+2}$ for some integer $j$ satisfying $1\leq j\leq e.$ Further, if $\mathfrak{S}_{e}(n)$ and $\mathfrak{U}_{e}(n)$ denote the total number of self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m},$ respectively, then by Lemma 2.1 and Remark 2.1, we get
@@
$\vskip-2.84526pt2^{\epsilon}(2^{m})^{\sum\limits_{i=1}^{s}\Lambda_{i}(n-\Lambda_{i+1})+\sum\limits_{j=1}^{s-1+\theta_{e}}\Lambda_{s+j}(n-\Lambda_{s+j+1}-\Lambda_{s+\theta_{e}-j})-\sum\limits_{a=1}^{s-\kappa_{1}-1}\Lambda_{a}-(1-\theta_{e})\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}+\mu}\hskip-5.69054pt\prod\limits_{\ell=s+1+\theta_{e}}^{e}{\lambda_{\ell}+n-\Lambda_{\ell}-\Lambda_{e+1-\ell}\brack\lambda_{\ell}}_{2^{m}}\vskip-2.84526pt$
@@
The desired result follows by applying Propositions 3.1, 3.3, 3.4, 3.6 and 3.8 when $e$ is even, whereas the desired result follows by applying Propositions 3.2, 3.3, 3.5, 3.6 and 3.8 when $e$ is odd.
@@
$\vskip-2.84526pt(2^{m})^{\sum\limits_{i=1}^{s}\Lambda_{i}(n-\Lambda_{i+1})+\sum\limits_{j=1}^{s-1+\theta_{e}}\Lambda_{s+j}(n-\Lambda_{s+j+1}-\Lambda_{s+\theta_{e}-j})-\sum\limits_{a=1}^{s-\kappa_{1}-1}\Lambda_{a}-(1-\theta_{e})\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}+\mu}\hskip-2.84526pt\prod\limits_{\ell=s+1+\theta_{e}}^{e}{\lambda_{\ell}+n-\Lambda_{\ell}-\Lambda_{e+1-\ell}\brack\lambda_{\ell}}_{2^{m}}\vskip-2.84526pt$
@@
When $e$ is even, the desired result follows by applying Propositions 3.1, 3.3, 3.7 and 3.8, while in the case when $e$ is odd, the desired result follows by applying Propositions 3.2, 3.3, 3.7 and 3.8. ∎
@@
    Now, for a given choice of $\mathcal{C}^{(s-\kappa_{1})},$ we will count the choices for a self-orthogonal code $\mathcal{C}^{(s+\theta_{e})}$ satisfying $\mathcal{C}^{(s-\kappa_{1})}\subseteq\mathcal{C}^{(s+\theta_{e})}\subseteq(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}.$ Towards this, we see, working as in case (I) in the proof of Theorem 3.1 of Yadav and Sharma [31], that $\mathcal{C}^{(s-\kappa_{1})}=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}}\rangle,$ where $\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}}$ are mutually orthogonal singular vectors in $\mathcal{I}(\mathcal{T}_{m}^{n})$ that are linearly independent over $\mathcal{T}_{m}.$ We further see, by [29, pp. 69-70], that the symplectic space $(\mathcal{I}(\mathcal{T}_{m}^{n}),B_{m}(\cdot,\cdot){\restriction_{\mathcal{I}(\mathcal{T}_{m}^{n})\times(\mathcal{I}(\mathcal{T}_{m}^{n})}})$ admits an orthogonal direct sum decomposition of the form: $\mathcal{I}(\mathcal{T}_{m}^{n})=\mathcal{X}\perp\mathcal{Y}$ with $\mathcal{X}=\langle\textbf{v}_{1},\textbf{v}_{1}^{\prime}\rangle\perp\langle\textbf{v}_{2},\textbf{v}_{2}^{\prime}\rangle\perp\cdots\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{v}_{\Lambda_{s-\kappa_{1}}}^{\prime}\rangle$ and $\mathcal{Y}=\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+1},\textbf{v}_{\Lambda_{s-\kappa_{1}}+1}^{\prime}\rangle\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+2},\textbf{v}_{\Lambda_{s-\kappa_{1}}+2}^{\prime}\rangle\perp\cdots\perp\langle\mathbf{v}_{\frac{n-1}{2}},\mathbf{v}_{\frac{n-1}{2}}^{\prime}\rangle,$ where $(\textbf{v}_{i},\textbf{v}_{i}^{\prime})$ is a hyperbolic pair in $\mathcal{I}(\mathcal{T}_{m}^{n})$ for $1\leq i\leq\frac{n-1}{2}.$ We further note that each $\textbf{w}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ can be uniquely written as $\textbf{w}=\sum\limits_{i=1}^{\frac{n-1}{2}}(\alpha_{i}\textbf{v}_{i}+\beta_{i}\textbf{v}_{i}^{\prime}),$ where $\alpha_{i},\beta_{i}\in\mathcal{T}_{m}$ for $1\leq i\leq\frac{n-1}{2}.$ As $\textbf{w}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}},$ we have $B_{m}(\textbf{w},\textbf{v}_{j})=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}},$ which implies that $\beta_{j}=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}}.$ Thus, each $\textbf{w}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ is of the form $\textbf{w}=\sum\limits_{i=1}^{\Lambda_{s-\kappa_{1}}}\alpha_{i}\textbf{v}_{i}+\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-1}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime}),$ which implies that $\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{w}\rangle=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-1}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime})\rangle.$ In view of this, we assume, without any loss of generality, that $\mathcal{C}^{(s+\theta_{e})}=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}}\rangle\perp\langle\textbf{w}_{1},\textbf{w}_{2},\ldots,\textbf{w}_{\Lambda_{s+\theta_{e}}-\Lambda_{s-\kappa_{1}}}\rangle,$ where $\textbf{w}_{1},\textbf{w}_{2},\ldots,\textbf{w}_{\Lambda_{s+\theta_{e}}-\Lambda_{s-\kappa_{1}}}$ are mutually orthogonal isotropic vectors in $\mathcal{Y}$ that are linearly independent over $\mathcal{T}_{m}.$ We further note that $\mathcal{Y}$ is a symplectic space over $\mathcal{T}_{m}$ of dimension $n-1-2\Lambda_{s-\kappa_{1}}$ and Witt index $\big(\frac{n-1}{2}\big)-\Lambda_{s-\kappa_{1}}.$ One can easily observe that the number of choices for a self-orthogonal code $\mathcal{C}^{(s+\theta_{e})}$ satisfying $\mathcal{C}^{(s-\kappa_{1})}\subseteq\mathcal{C}^{(s+\theta_{e})}\subseteq(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ is equal to the number of $(\Lambda_{s+\theta_{e}}-\Lambda_{s-\kappa_{1}})$-dimensional totally isotropic subspaces of the symplectic space $\mathcal{Y},$ which, by Exercise 8.1 (ii) of [29], has precisely $\prod\limits_{i=\Lambda_{s-\kappa_{1}}}^{\Lambda_{s+\theta_{e}}-1}\left(\frac{2^{m(n-2i-1)}-1}{2^{m(i+1-\Lambda_{s-\kappa_{1}})}-1}\right)$ distinct choices. Finally, for a given choice of $\mathcal{C}^{(s-\kappa_{1})},$ we note that the chain $\mathcal{C}^{(s-\kappa_{1}+1)}\subseteq\mathcal{C}^{(s-\kappa_{1}+2)}\subseteq\cdots\subseteq\mathcal{C}^{(s+\theta_{e})}$ of codes has precisely
@@
      Next, if $\Lambda_{s+\theta_{e}}\neq\Lambda_{s-\kappa_{1}},$ for a given choice of $\mathcal{C}^{(s-\kappa_{1})},$ we need to count the choices for a self-orthogonal code $\mathcal{C}^{(s+\theta_{e})}$ satisfying $\mathcal{C}^{(s-\kappa_{1})}\subseteq\mathcal{C}^{(s+\theta_{e})}\subseteq(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}.$ Here, we see, by [29, pp. 69-70], that the space $(\mathcal{I}(\mathcal{T}_{m}^{n}),B_{m}(\cdot,\cdot){\restriction_{\mathcal{I}(\mathcal{T}_{m}^{n})\times(\mathcal{I}(\mathcal{T}_{m}^{n})}})$ admits an orthogonal direct sum decomposition of the form: $\mathcal{I}(\mathcal{T}_{m}^{n})=\mathcal{X}_{0}\perp\mathcal{Y}_{0}\perp\langle\textbf{1}\rangle$ with $\mathcal{X}_{0}=\langle\textbf{v}_{1},\textbf{v}_{1}^{\prime}\rangle\perp\langle\textbf{v}_{2},\textbf{v}_{2}^{\prime}\rangle\perp\cdots\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{v}_{\Lambda_{s-\kappa_{1}}}^{\prime}\rangle$ and $\mathcal{Y}_{0}=\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+1},\textbf{v}_{\Lambda_{s-\kappa_{1}}+1}^{\prime}\rangle\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+2},\textbf{v}_{\Lambda_{s-\kappa_{1}}+2}^{\prime}\rangle\perp\cdots\perp\langle\mathbf{v}_{\frac{n-2}{2}},\mathbf{v}_{\frac{n-2}{2}}^{\prime}\rangle,$ where $(\textbf{v}_{i},\textbf{v}_{i}^{\prime})$ is a hyperbolic pair in $\mathcal{I}(\mathcal{T}_{m}^{n})$ for $1\leq i\leq\frac{n-2}{2}.$ We further note that each $\textbf{z}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ can be uniquely written as $\textbf{z}=\sum\limits_{i=1}^{\frac{n-2}{2}}(\alpha_{i}\textbf{v}_{i}+\beta_{i}\textbf{v}_{i}^{\prime})+\gamma\textbf{1},$ where $\alpha_{i},\beta_{i},\gamma\in\mathcal{T}_{m}$ for $1\leq i\leq\frac{n-2}{2}.$ Since $\textbf{z}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}},$ we have $B_{m}(\textbf{z},\textbf{v}_{j})=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}},$ which implies that $\beta_{j}=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}}.$ Thus, each $\textbf{z}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ is of the form $\textbf{z}=\sum\limits_{i=1}^{\Lambda_{s-\kappa_{1}}}\alpha_{i}\textbf{v}_{i}+\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-2}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime})+\gamma\textbf{1},$ which implies that $\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{z}\rangle=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-2}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime})+\gamma\textbf{1}\rangle.$ In view of this, we assume, without any loss of generality, that the code $\mathcal{C}^{(s+\theta_{e})}$ is of the following two forms:
@@
When $2\kappa\leq e,$ the desired results follows by Lemmas 4.1, 4.3, 4.4 and 4.5. On the other hand, when $2\kappa>e,$ we get the desired result by Lemmas 4.2, 4.3, 4.4 and 4.6. ∎

### v2

@@ 4 Enumeration formulae for self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m}$
In this section, we will obtain explicit enumeration formulae for self-orthogonal and self-dual codes of an arbitrary length over $\mathscr{R}_{e,m}.$ In particular, when $e=2,$ one can easily see that either $\mathscr{R}_{e,m}\simeq\mathbb{F}_{2^{m}}[u]/\langle u^{2}\rangle$ or $\mathscr{R}_{e,m}\simeq GR(4,m).$ The enumeration formula for self-dual codes over $\mathbb{F}_{2^{m}}[u]/\langle u^{2}\rangle$ was derived by Gaborit [15, Th. 3], while the enumeration formula for self-orthogonal codes over $\mathbb{F}_{2^{m}}[u]/\langle u^{2}\rangle$ was obtained by Galvez et al. [17, Th. 2]. Yadav and Sharma [34, Th. 5.1] counted all self-orthogonal and self-dual codes of an arbitrary length over $GR(4,m).$ In view of this, we assume, throughout this section, that $e\geq 3.$ We also assume that $n$ is a positive integer and $\lambda_{1},\lambda_{2},\ldots,\lambda_{e+1}$ are non-negative integers satisfying $n=\lambda_{1}+\lambda_{2}+\cdots+\lambda_{e+1}.$ Now, let $\mathfrak{S}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})$ and $\mathfrak{U}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})$ denote the number of distinct self-orthogonal and self-dual codes of type $\{\lambda_{1},\lambda_{2},\ldots,\lambda_{e}\}$ and length $n$ over $\mathscr{R}_{e,m},$ respectively. We note, by Lemma 2.1 and Remark 2.1, that $\mathfrak{S}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})=0$ if $2\lambda_{1}+2\lambda_{2}+\cdots+2\lambda_{e-j+1}+\lambda_{e-j+2}+\cdots+\lambda_{j}>n$ holds for some integer $j$ satisfying $s+1\leq j\leq e,$ while $\mathfrak{U}_{e}(n;\lambda_{1},\lambda_{2},\ldots,\lambda_{e})=0$ if $\lambda_{j}\neq\lambda_{e-j+2}$ for some integer $j$ satisfying $1\leq j\leq e.$ Further, if $\mathfrak{S}_{e}(n)$ and $\mathfrak{U}_{e}(n)$ denote the total number of self-orthogonal and self-dual codes of length $n$ over $\mathscr{R}_{e,m},$ respectively, then by Lemma 2.1 and Remark 2.1, we get
@@
$2^{\epsilon}(2^{m})^{\sum\limits_{i=1}^{s}\Lambda_{i}(n-\Lambda_{i+1})+\sum\limits_{j=1}^{s-1+\theta_{e}}\Lambda_{s+j}(n-\Lambda_{s+j+1}-\Lambda_{s+\theta_{e}-j})-\sum\limits_{a=1}^{s-\kappa_{1}-1}\Lambda_{a}-(1-\theta_{e})\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}+\mu}\hskip-5.69054pt\prod\limits_{\ell=s+1+\theta_{e}}^{e}{\lambda_{\ell}+n-\Lambda_{\ell}-\Lambda_{e+1-\ell}\brack\lambda_{\ell}}_{2^{m}}\vskip-2.84526pt$
@@
The desired result follows by applying Propositions 3.1 and 3.33.5 when $e$ is even, whereas the desired result follows by applying Propositions 3.23.5 when $e$ is odd.
@@
$(2^{m})^{\sum\limits_{i=1}^{s}\Lambda_{i}(n-\Lambda_{i+1})+\sum\limits_{j=1}^{s-1+\theta_{e}}\Lambda_{s+j}(n-\Lambda_{s+j+1}-\Lambda_{s+\theta_{e}-j})-\sum\limits_{a=1}^{s-\kappa_{1}-1}\Lambda_{a}-(1-\theta_{e})\frac{\Lambda_{s}(\Lambda_{s}-1)}{2}+\mu}\hskip-2.84526pt\prod\limits_{\ell=s+1+\theta_{e}}^{e}{\lambda_{\ell}+n-\Lambda_{\ell}-\Lambda_{e+1-\ell}\brack\lambda_{\ell}}_{2^{m}}\vskip-2.84526pt$
@@
When $e$ is even, the desired result follows by applying Propositions 3.1, 3.3 and 3.5, while in the case when $e$ is odd, the desired result follows by applying Propositions 3.2, 3.3 and 3.5. ∎
@@
    Now, for a given choice of $\mathcal{C}^{(s-\kappa_{1})},$ we will count the choices for a self-orthogonal code $\mathcal{C}^{(s+\theta_{e})}$ satisfying $\mathcal{C}^{(s-\kappa_{1})}\subseteq\mathcal{C}^{(s+\theta_{e})}\subseteq(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}.$ Towards this, we see, working as in case (I) in the proof of Theorem 3.1 of Yadav and Sharma [34], that $\mathcal{C}^{(s-\kappa_{1})}=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}}\rangle,$ where $\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}}$ are mutually orthogonal singular vectors in $\mathcal{I}(\mathcal{T}_{m}^{n})$ that are linearly independent over $\mathcal{T}_{m}.$ We further see, by [32, pp. 69-70], that the symplectic space $(\mathcal{I}(\mathcal{T}_{m}^{n}),B_{m}(\cdot,\cdot){\restriction_{\mathcal{I}(\mathcal{T}_{m}^{n})\times(\mathcal{I}(\mathcal{T}_{m}^{n})}})$ admits an orthogonal direct sum decomposition of the form: $\mathcal{I}(\mathcal{T}_{m}^{n})=\mathcal{X}\perp\mathcal{Y}$ with $\mathcal{X}=\langle\textbf{v}_{1},\textbf{v}_{1}^{\prime}\rangle\perp\langle\textbf{v}_{2},\textbf{v}_{2}^{\prime}\rangle\perp\cdots\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{v}_{\Lambda_{s-\kappa_{1}}}^{\prime}\rangle$ and $\mathcal{Y}=\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+1},\textbf{v}_{\Lambda_{s-\kappa_{1}}+1}^{\prime}\rangle\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+2},\textbf{v}_{\Lambda_{s-\kappa_{1}}+2}^{\prime}\rangle\perp\cdots\perp\langle\mathbf{v}_{\frac{n-1}{2}},\mathbf{v}_{\frac{n-1}{2}}^{\prime}\rangle,$ where $(\textbf{v}_{i},\textbf{v}_{i}^{\prime})$ is a hyperbolic pair in $\mathcal{I}(\mathcal{T}_{m}^{n})$ for $1\leq i\leq\frac{n-1}{2}.$ We further note that each $\textbf{w}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ can be uniquely written as $\textbf{w}=\sum\limits_{i=1}^{\frac{n-1}{2}}(\alpha_{i}\textbf{v}_{i}+\beta_{i}\textbf{v}_{i}^{\prime}),$ where $\alpha_{i},\beta_{i}\in\mathcal{T}_{m}$ for $1\leq i\leq\frac{n-1}{2}.$ As $\textbf{w}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}},$ we have $B_{m}(\textbf{w},\textbf{v}_{j})=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}},$ which implies that $\beta_{j}=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}}.$ Thus, each $\textbf{w}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ is of the form $\textbf{w}=\sum\limits_{i=1}^{\Lambda_{s-\kappa_{1}}}\alpha_{i}\textbf{v}_{i}+\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-1}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime}),$ which implies that $\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{w}\rangle=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-1}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime})\rangle.$ In view of this, we assume, without any loss of generality, that $\mathcal{C}^{(s+\theta_{e})}=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}}\rangle\perp\langle\textbf{w}_{1},\textbf{w}_{2},\ldots,\textbf{w}_{\Lambda_{s+\theta_{e}}-\Lambda_{s-\kappa_{1}}}\rangle,$ where $\textbf{w}_{1},\textbf{w}_{2},\ldots,\textbf{w}_{\Lambda_{s+\theta_{e}}-\Lambda_{s-\kappa_{1}}}$ are mutually orthogonal isotropic vectors in $\mathcal{Y}$ that are linearly independent over $\mathcal{T}_{m}.$ We further note that $\mathcal{Y}$ is a symplectic space over $\mathcal{T}_{m}$ of dimension $n-1-2\Lambda_{s-\kappa_{1}}$ and Witt index $\big(\frac{n-1}{2}\big)-\Lambda_{s-\kappa_{1}}.$ One can easily observe that the number of choices for a self-orthogonal code $\mathcal{C}^{(s+\theta_{e})}$ satisfying $\mathcal{C}^{(s-\kappa_{1})}\subseteq\mathcal{C}^{(s+\theta_{e})}\subseteq(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ is equal to the number of $(\Lambda_{s+\theta_{e}}-\Lambda_{s-\kappa_{1}})$-dimensional totally isotropic subspaces of the symplectic space $\mathcal{Y},$ which, by Exercise 8.1 (ii) of [32], has precisely $\prod\limits_{i=\Lambda_{s-\kappa_{1}}}^{\Lambda_{s+\theta_{e}}-1}\left(\frac{2^{m(n-2i-1)}-1}{2^{m(i+1-\Lambda_{s-\kappa_{1}})}-1}\right)$ distinct choices. Finally, for a given choice of $\mathcal{C}^{(s-\kappa_{1})},$ we note that the chain $\mathcal{C}^{(s-\kappa_{1}+1)}\subseteq\mathcal{C}^{(s-\kappa_{1}+2)}\subseteq\cdots\subseteq\mathcal{C}^{(s+\theta_{e})}$ of codes has precisely
@@
      Next, if $\Lambda_{s+\theta_{e}}\neq\Lambda_{s-\kappa_{1}},$ for a given choice of $\mathcal{C}^{(s-\kappa_{1})},$ we need to count the choices for a self-orthogonal code $\mathcal{C}^{(s+\theta_{e})}$ satisfying $\mathcal{C}^{(s-\kappa_{1})}\subseteq\mathcal{C}^{(s+\theta_{e})}\subseteq(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}.$ Here, we see, by [32, pp. 69-70], that the space $(\mathcal{I}(\mathcal{T}_{m}^{n}),B_{m}(\cdot,\cdot){\restriction_{\mathcal{I}(\mathcal{T}_{m}^{n})\times(\mathcal{I}(\mathcal{T}_{m}^{n})}})$ admits an orthogonal direct sum decomposition of the form: $\mathcal{I}(\mathcal{T}_{m}^{n})=\mathcal{X}_{0}\perp\mathcal{Y}_{0}\perp\langle\textbf{1}\rangle$ with $\mathcal{X}_{0}=\langle\textbf{v}_{1},\textbf{v}_{1}^{\prime}\rangle\perp\langle\textbf{v}_{2},\textbf{v}_{2}^{\prime}\rangle\perp\cdots\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{v}_{\Lambda_{s-\kappa_{1}}}^{\prime}\rangle$ and $\mathcal{Y}_{0}=\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+1},\textbf{v}_{\Lambda_{s-\kappa_{1}}+1}^{\prime}\rangle\perp\langle\textbf{v}_{\Lambda_{s-\kappa_{1}}+2},\textbf{v}_{\Lambda_{s-\kappa_{1}}+2}^{\prime}\rangle\perp\cdots\perp\langle\mathbf{v}_{\frac{n-2}{2}},\mathbf{v}_{\frac{n-2}{2}}^{\prime}\rangle,$ where $(\textbf{v}_{i},\textbf{v}_{i}^{\prime})$ is a hyperbolic pair in $\mathcal{I}(\mathcal{T}_{m}^{n})$ for $1\leq i\leq\frac{n-2}{2}.$ We further note that each $\textbf{z}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ can be uniquely written as $\textbf{z}=\sum\limits_{i=1}^{\frac{n-2}{2}}(\alpha_{i}\textbf{v}_{i}+\beta_{i}\textbf{v}_{i}^{\prime})+\gamma\textbf{1},$ where $\alpha_{i},\beta_{i},\gamma\in\mathcal{T}_{m}$ for $1\leq i\leq\frac{n-2}{2}.$ Since $\textbf{z}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}},$ we have $B_{m}(\textbf{z},\textbf{v}_{j})=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}},$ which implies that $\beta_{j}=0$ for $1\leq j\leq\Lambda_{s-\kappa_{1}}.$ Thus, each $\textbf{z}\in(\mathcal{C}^{(s-\kappa_{1})})^{\perp_{B_{m}}}$ is of the form $\textbf{z}=\sum\limits_{i=1}^{\Lambda_{s-\kappa_{1}}}\alpha_{i}\textbf{v}_{i}+\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-2}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime})+\gamma\textbf{1},$ which implies that $\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\textbf{z}\rangle=\langle\textbf{v}_{1},\textbf{v}_{2},\ldots,\textbf{v}_{\Lambda_{s-\kappa_{1}}},\sum\limits_{j=\Lambda_{s-\kappa_{1}}+1}^{\frac{n-2}{2}}(\alpha_{j}\textbf{v}_{j}+\beta_{j}\textbf{v}_{j}^{\prime})+\gamma\textbf{1}\rangle.$ In view of this, we assume, without any loss of generality, that the code $\mathcal{C}^{(s+\theta_{e})}$ is of the following two forms:
@@
When $2\kappa\leq e,$ the desired result follows by Lemmas 4.1, 4.3, 4.4 and 4.5. On the other hand, when $2\kappa>e,$ we get the desired result by Lemmas 4.2, 4.3, 4.4 and 4.6. ∎